6/15/99 evening shift

Yee Bob Hsiung (hsiung@fnal.gov)
Wed, 16 Jun 1999 00:26:50 -0500 (CDT)

Shift Summary for evening shift, 6/15/99

On Shift: Bob Hsiung, Greg Bock

Hours of beam delivered at all intensities: 7.5 with lots abort
Hours of beam delivered at nominal intensity: 5
Hours of E832 data (special runs not included): 7 with lots pipelineerrors
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:
Lots init pipelines and kfrends, and run changes.
Lost permit for stream 5 a few times. Cycle repeater power on/off once.

========================================================================
Runs ended this shift:
Run 13611 - 0.71M events - BEAM_832 DC1,2 HV=2400, del_V=100V for window
Run 13612 - 1.06M events - BEAM_832 back to del_V=0
Run 13613 - 0.95M events - BEAM_832 "
Run 13614 - 2.47M events - BEAM_832 DC1 HV=2350V del_V=0
Run 13615 - in progress, - BEAM_832 DC1,2 HV=2400, del_V=100V,
targeting changed 3/4mm west
Tapes brought to FCC this shift: None
(including tapes for online split)

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:
Startup after today's downtime. MCR tuning for intensity.
Beam were sporadic for the first few hours. After intensity raised
to 4E12, beam is stable but with many aborts. Lots pipeline errors
at 4E12.
Neal Lane (President's science advisor) visits KTeV.
Drift chamber ineff vs HV study by chamber group, took runs
with window wire HV 100V lower than cathode wire HV for DC1 & 2
and a few other conditions.
Targeting study by RickFord.
We can get 3500 KL->2pi0 in 2 hours at 4E12 (=E731A statistics).

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13611 15:35 startup waiting for beam, beam stop closed
initially. DC window HV is lowered by 100V.
16:00 Neal Lane visits control room. still no beam.
16:22 low intensity about 1E12, 5 booster batches.
MCR tuning for losses.
13612 17:50 DC HV back to nominal. Added new B09 for MA
mip monitor.
19:40 intensity up to 3E12, 11 booster batches, BOA=1.7
13613 19:50 4E12 intensity run,
Lots pipeline error.
13614 20:57 lower DC1 HV=2350.

13615 23:16 Repeat DC HV setting as run 13611 at 4E12.
Rick Ford changed the beam targeting 3/4 mm west.
Rick K bumps up the accidentals by x4.

========================================================================
Follow up Notes on Previous Shift

Crate 16 is OK now after day shift replaced it.

========================================================================
Items to Follow up on Next Shift

At 4E12 intensity (BOA=1.5), we have frequent pipeline errors which
requires init pipeline and kfrend almost every 5-10 min.

One of the trim magnet NM2H does not work for targeting study, so the
targeting angle study by Rick Ford will have to wait till tomorrow.

Contiue drift chamber ineff vs HV (Cathode and Window) study.

If there is a long break of beam, take a long laser scan for
Peter Shanahan.